OpenStudy (anonymous):

solve for t : t^6+4t^3=45

OpenStudy (akshay_budhkar):

let\[t^3=u\] it converts into an easy solving quadratic equation

OpenStudy (anonymous):

start with \[t^6+4t^3-45=0\] then factor. if it is not clear how to factor write as \[x^2+4x-45=0\] and factor that

OpenStudy (akshay_budhkar):

\[u^2+4u-45=0\]

OpenStudy (akshay_budhkar):

Can u do from here?

OpenStudy (anonymous):

can you help still from this point?

OpenStudy (akshay_budhkar):

sure\[\begin{array} \\u^2+9u-5u-45=0 \\ \therefore u(u+9)-5(u+9)=0 \\ \therefore (u-5)(u+9)=0\\ u=5 or u= -9 \\ \therefore t^3=5 or t^3=-9 \end {array}\]

OpenStudy (anonymous):

thank you!!! this helped a lot!

OpenStudy (akshay_budhkar):

Yw :D

OpenStudy (anonymous):

wait so is the answer 5 or -9 or is it something else too?